OpenStudy (anonymous):

Log (x) + Log (x-3) = Log (3x) Please tell me whats the procedure for this eq.

OpenStudy (anonymous):

Log (x) + Log (x-3) = Log (3x) Log( x(x-3)) =Log (3 x) x(x-3) =3x

OpenStudy (anonymous):

\[ x^2-3 x =3x\\ x^2-6x=0\\ x=0 \\ x=6\\ \] x=0 is not acceptable. Why?

OpenStudy (anonymous):

is it x=0 and x=6?

OpenStudy (anonymous):

x= 0 is not good since Log[0] is not defined.

OpenStudy (anonymous):

why? is it because it has to be x>1?

OpenStudy (anonymous):

The answer is x=6

OpenStudy (anonymous):

okay thanks, i didnt know about the 0 Rule, thats why i was wrong :D thank you :D

OpenStudy (anonymous):

yw .

OpenStudy (anonymous):

sorry, i was looking for your Profile :)